OpenStudy (anonymous):

Let f(x) = x + 2 and g(x) = x2 Find ( g o f)(-5)

OpenStudy (aum):

( g o f)(-5) = g( f(-5) ) f(x) = x + 2 f(-5) = -5 + 2 = -3 g(x) = x^2 g( f(-5) ) = g(-3) = (-3)^2 = 9

OpenStudy (isaiah.feynman):